The waves on a string have wave speed 8.00m/s, amplitude 0.0600m, and wavelength 0.400m. The waves travel in the -x direction and at t = 0s, the end of the string has its maximum upward displacement.


(a) Write a wave function describing the wave and


(b). Find the transverse displacement of a particle at x = 0.350m at time t = 0.150s

1
Expert's answer
2022-04-11T12:44:51-0400

(a) Write a wave function describing the wave


"y(x,t)=A\cos(2\pi ft+\frac{2\pi}{\lambda}x)""f=v/\lambda=8.0/0.400=20\:\rm Hz""y(x,t)=0.0600{\:\rm m}\cos(40\pi t+5\pi x)"

(b) Find the transverse displacement of a particle at x = 0.350m at time t = 0.150s


"y=0.0600{\:\rm m}\cos(40\pi *0.150+5\pi *0.350)\\
=0.0424\:\rm m"